Grade 11Physical Chemistry

A bottle of whisky contains 12%aqueous ethanol by volume. The density of ethanol (C2H5OH) is
0.8 g/cm3 .Calculate.

HARSH2OP
5 Years agoGrade 11

1 Answer

Ayush Raj
5 Years ago
Given %v/v =12%, means 12ml solute(ethanol) in 100ml solution.
 
Thus, mass of ethanol =v×d =12×0.8 =9.6gm
       moles of ethanol=9.6/46 mol
 
   Volume of solvent =(100-12)ml=88ml
  Here, solvent H2O , so density of solvent =1g/ml
      Mass of solvent =88×1=88gm
  
 Mass of solution=mass of solute+mass of solvent
                            =9.6+88=97.6gm
 
Now, a) molality = moles of solute /mass of solvent in kg
   =9.6/46 ÷88/1000 = 2.37m
 
       b) Molarity =moles of solute/volume of solution in Litre 
     =9.6/46 ÷ 100/1000 = 2.08M
 
   c) %w/w =(mass of solute/mass of solution) ×100
                =(9.6 ÷ 97.6) ×100 = 9.83%
 
Thanks. 
It was a brilliant question, where most important thing was to take volume conserved.